OpenStudy (anonymous):

The slope of the line whose equation is 5x - 3y = 4 is 5/3??

OpenStudy (anonymous):

If you put it in slope intercept form it makes it easy to see. You're correct! :)

OpenStudy (anonymous):

-3y = -5x + 4 3y = 5x - 4 y = 5/3(x) - (4/3) Slope = 5/3

OpenStudy (amistre64):

ax + by = c slope = -a/b .............................. 5x -3y = 4 slope = -5/-3 = 5/3 ; yep